OpenStudy (anonymous):

Change 2x + 5y = –30 to slope-intercept form. You must show all work.

OpenStudy (anonymous):

2x+5y=-30 5y=-2-30 y=-2/5x-6

OpenStudy (anonymous):

5y=-2x-30 for that second line